OpenStudy (anonymous):

The instructions for building a kite tell you to begin by cutting a strip of wood to 45 cm in length, but your ruler only measures inches. How many inches should the piece of wood be? Round to the nearest quarter of an inch

OpenStudy (anonymous):

@MathLegend

OpenStudy (anonymous):

2.54cm=1 inch 1 cm=1/2.54 inch 45 cm=1/2.54 *54 inch=?